OpenStudy (anonymous):

Simple math question: Can my surface area change if my volume stays the same?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

The volume of a sphere = 4/3 * pi * r^3 The volume of a cube has = x^3 where x = length of an edge. if you set these to be = lets say 1cm^3 as teh area and solve for x and r each

OpenStudy (anonymous):

we have 4/3*pi*r^3 = 1 means r ~ 0.62 and we have x^3 = 1 means x = 1

OpenStudy (anonymous):

the surface area of a cube = 6 * x^2 = 6 The surface area of a sphere = 4pi * r^2 = 4.83

OpenStudy (anonymous):

So my problem says, a can has the volume of 200, and i am to find the dimensions of said can, so that they are the most cost effective (smallest)

OpenStudy (anonymous):

you want to minimize surface area while maximizing volume.

OpenStudy (anonymous):

yes but the volume must stay at 200

OpenStudy (anonymous):

V = pi*r^2 * h SA = 2pi*rh + 2pi*r^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the dimensions must be altered so i find the ideal surface area

OpenStudy (anonymous):

One second Let me work on it on paper first.

OpenStudy (anonymous):

What math are you in algebra or calc?

OpenStudy (anonymous):

ok, i was thinking of just plugging in functions until i find the answer out.

OpenStudy (anonymous):

If you solve volume for h, you get h = 200/(pi*r^2)

OpenStudy (anonymous):

plug this into the h in the SA equation to get : \[SA = 2\pi r(200/(\pi*r^2))+ 2\pi*r\]

OpenStudy (anonymous):

Now, what you want to do is solve this equation for a value of r that gives the smallest SA

OpenStudy (anonymous):

so which equation a i solving first??

OpenStudy (anonymous):

you solve volume for the variable h. then you plug in what h equals into the h in the surface area equation

OpenStudy (anonymous):

ohhhhhhh makes sense

OpenStudy (anonymous):

wait, so do i plug in 200 as SA?? @manumben

OpenStudy (anonymous):

Which math class are you in? calculus?

OpenStudy (anonymous):

Honors Geometry

OpenStudy (anonymous):

ok then im helping you the wrong way... one moment

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

h = 2r is the most efficient dimensions for a can

OpenStudy (anonymous):

this is because it is closest to a sphere which has the greatest surface area to volume ratio

OpenStudy (anonymous):

200 = pi * r^2 * h substitute using h= 2r or r = 1/2 h

OpenStudy (anonymous):

so then as long as my height is twice as big as my length and the volume multiplies to 200 im good

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Noice!

OpenStudy (anonymous):

but the surface area of a 10X20 can is 1668, so wouldnt it be more efficient to use a can with a radius of 5 and height of 3??

OpenStudy (anonymous):

@manumben

OpenStudy (anonymous):

@manumben